# Large Language Models Reflect Human Citation Patterns with a Heightened Citation Bias

### v2

@@ 6 Discussion
We designed an experiment to explore the intrinsic citation behavior of GPT-4 and its potential biases when generating scholarly references. While previous work focuses on LLMs generating short papers or literature reviews [27, 35], we let GPT-4 generate suggestions of scholarly references for anonymized in-text citations. Importantly, we do not enhance the LLM through search and retrieval-augmented generation, but evaluate the model’s internalization of citation patterns in its parametric knowledge encapsulated during its (pre-)training.
@@
Our findings are significant as they represent a first step towards understanding the real-world impact of LLMs in scientific research [3, 44]. By highlighting the heightened citation bias in GPT-4 generated references, we demonstrate the model’s tendency to favor highly cited papers, which could exacerbate existing biases in scientific discourse. This evaluation moves beyond more traditional LLM benchmarks [30], emphasizing the practical implications of deploying these models in academic contexts [12]. The results suggest that while LLMs have the potential to streamline various aspects of research, careful consideration is needed to mitigate the amplification of biases.
@@
One plausible hypothesis for the heightened citation bias observed in GPT-4 is the increased frequency of citations to heavily cited papers within the model’s training data. This prevalence makes these references more likely to be generated accurately and recognized as existent. Additionally, such biases may stem from generic training effects, where models preferentially learn patterns that are more common in the data, leading to biases towards shorter titles, more heavily cited, and slightly less recent works. These tendencies may persist despite improvements in data quantity or model sophistication [15, 19].
@@
We developed and open-source an extensible, automated pipeline to systematically analyze the references generated by LLMs. Although our methodology is robust, it is not without limitations. The use of simple prompts and the zero-shot setting [17] aims to minimize bias in the generation process, but this simplicity might not capture the full spectrum of potential LLM capabilities. There are numerous alternative approaches and prompt designs that future research can explore to enhance the accuracy and relevance of generated references [37, 39, 42]. However, our iterative approach already indicates that biases remain inherent in these generations.
@@
In conclusion, while LLMs like GPT-4 can significantly aid in citation generation, they also risk amplifying existing biases and introducing new ones, potentially skewing the structuring and the dissemination of scientific knowledge. Our study underscores the necessity for developing balanced methods to interact with LLMs, incorporating diverse datasets, and implementing bias mitigation strategies. Fair prompting techniques [22], for instance, can be employed to reduce bias, but continuous vigilance and methodological innovation are required to ensure that the integration of LLMs into academic workflows promotes accurate knowledge dissemination.
Acknowledgments and Disclosure of Funding
@@
Appendix A Data
We describe the steps of our automated pipeline to retrieve all the necessary information for our analysis. Our data collection resulted in $166$ papers published at AAAI ($25$), NeurIPS ($72$), ICML ($38$), and ICLR ($31$) for a total of 3,066 references (see Appendix Table LABEL:tab:paper_details for a full list of included papers). The data collection pipeline uses GPT-4 to postprocess parts of the data, which costs approximately 14 dollars for our experiment. Note that these steps only have to be carried out once for the data collection. However, steps 4 and 5 are also used to postprocess and enrich the information of the generated references and will need to be carried out for each run. The experiment was run on 4 November 2023 and each step was manually verified and tested.
@@ Appendix A Data
We search for all papers on arXiv originally posted between 1 March 2022 and 31 October 2023 in the machine learning (cs.LG) category which refer to AAAI, NeurIPS, ICLR, or ICML in their journal reference. Note that we also verify whether we can use all these arXiv papers given their data licenses and attribute their participation in Appendix Table LABEL:tab:paper_details. We use keywords (i.e., workshop, tiny paper, 2020, 2021, track on datasets and benchmarks, and bridge) to remove papers that do not appear in the conference proceedings or earlier than 2022. We download and unzip the tar.gz file provided by the authors to arXiv and check whether the paper exists on Semantic Scholar via title matching. We store the title, ID, and date from arXiv and Semantic Scholar. Additionally, we store all the reference titles with their corresponding ID from Semantic Scholar [16].
@@
Below, we share with you a list of references with their corresponding citation number between square brackets. Could you for each reference extract the authors, the number of authors, title, publication year, and publication venue? Please only return the extracted information in a markdown table with the citation number (without brackets), authors, number of authors, title, publication year, and publication venue as columns. $=$ [GPT-4 generated reference list]
@@
We enrich the information from the introduction references by matching the extracted title from the csv file in the previous step with the reference titles that we extracted from Semantic Scholar in step 1. This approach provides an extra check that GPT-4 does not change the title information in Step 4. After matching, we can use the Semantic Scholar ID to retrieve the publication venue, year, authors, citation count, influential citation count, and reference count [16]. Additionally, we store the IDs of the papers to which the introduction references themselves refer.
@@
We prompt GPT-4 with the main content, which includes the author information, conference information, abstract, and introduction, accompanied by the default system message: “You are a helpful assistant” and the following prompt:
@@
We determine whether the generated references exist via title and author names matching with Semantic Scholar entries [16]. We search Semantic Scholar for the three best matches based on the reference’s title and then compute the title and author names similarity. For titles, we measure the similarity between the Semantic Scholar match and the generated reference by comparing the best matching substring. For authors, we compare them by splitting into tokens (words), removing duplicates, and then calculating the similarity based on the best partial match of the sets of tokens. In case of “et al.,” we only consider the first author. The similarity is computed by character-level comparison. We determined the thresholds for the title and authors scores by manually labelling $100$ matches as true or false and minimizing the false positive rate. We obtain on this sample an accuracy of $95\%$ with $5$ false positives, i.e. generated references falsely classified as non-existent.
@@
We also build on our “vanilla” approach, by introducing an “iterative” approach where we prompt GPT-4 with the main content accompanied by the default system message: “You are a helpful assistant” and the following prompt:
[vanilla prompt + GPT-4’s response ] The following references associated with these citation numbers: [numbers of non-existent generated references ] do not exist. Can you replace all these non-existent references with existing ones? Keep the other references as they are. Do not return anything except the citation number between square brackets and the corresponding reference. $=$ [main content]
@@ Appendix B Additional Figures
Figure B1: Properties of the ground truth and generated introduction references for the iterative strategy are consistent with the properties of the vanilla strategy. This figure displays the properties of the ground truth ($n=5,178$, in blue) and generated references ($n=5,178$, in green), further subdividing the generated references into existing ($n=3,244$, in orange) and non-existent categories ($n=1,934$, in red), from the original data sources of five runs for the iterative strategy with GPT-4. Note that these are the references which are labelled “non-existent” in the vanilla strategy. a, b, c, d, e, f and g, The iterative results exhibit very similar properties to the vanilla results shown in Fig. 1.
@@
Figure B2: The properties of the existing generated references and their corresponding ground truth are consistent between the original data sources and Semantic Scholar data. This figure compares the computation of the properties of the existing generated references ($n=9.376$) and their corresponding ground truth ($n=9.376$) between the data from the original sources (in dark blue and orange, as shown in Fig. 1) and from Semantic Scholar (in light blue and orange). a, The distributions of the number of characters in title for the existing generated references and their corresponding ground truth are very similar between the data from the original sources and Semantic Scholar. b, There is a discrepancy between the data from the original sources and Semantic Scholar for the number of authors in the existing generated references due to the extensive use of “et al”. This discrepancy results in a relatively larger portion of three authors or more, but does not change the previous conclusions. c, The distributions over time are very similar between the data from the original sources and Semantic Scholar. d, There is a discrepancy between the data from the original sources and Semantic Scholar for the publication venues. The discrepancy is consistent across the existing generated references and their corresponding ground truth as both have a lower number of arXiv papers and a larger number of ICLR papers.
Figure B3: Ground truth papers which correspond to existing generated references that appear in the paper have substantially more citations. This figure displays the properties of the ground truth references which correspond to existing generated references ($n=9,376$, in yellow), the subset of existing generated references which appear in the paper itself ($n=2,474$, in blue), and the non-existent generated references ($n=5,178$, in green), from the original data sources of five runs for the vanilla strategy with GPT-4. a, b, c, d, e and f, The ground truth papers which correspond to existing references which appear in the paper have by far the most citations, followed by the existing references, and the ground truth papers corresponding to non-existent references have the lowest numbers of citations. These findings further indicate the tendency for LLMs to more easily generate references to highly cited papers. The distributions of all other characteristics are very similar.
@@
Figure B4: A high consistency in publication venue distributions between ground truth and GPT-4 generated references with a notable bias towards arXiv, NeurIPS, and “Others”. This figure displays the distributions and pairwise transition of publication venues for ground truth and generated references at the conference and individual paper and reference level. a and b, We observe that the distributions of publication venues for both ground truth and generated references are very similar across the various conferences, i.e., AAAI, NeurIPS, ICML, and ICLR. c and d, The pairwise transition matrix from ground truth to generated publication venues at the reference level indicates a large overall agreement, but with a strong preference in GPT-4 generated references for arXiv, NeurIPS, and “Others” in the case of disagreement. The preference for NeurIPS may be due to the relatively large number of NeurIPS papers in our sample and the large share of arXiv and “Others” points to favoring a wider array of venues which may potentially dilute the perceived relevance of key conferences. e, The scatter plot shows for each paper the number of ground truth references to one of the top conferences, i.e., AAAI, NeurIPS, ICML, and ICLR, and the corresponding number of generated references (×$\times 5$ for five runs) which refer to the same conference.The strong pairwise correlation between the ground truth and generated references to the top conferences at the individual paper level affirms the high consistency in publication venue distributions between ground truth and GPT-4 generated references.
Figure B5: The influential citation bias in existing generated references is unrelated to the recency of ground truth references. This figure shows that the existing generated references ($n=9,376$, in orange) consistently exhibit a higher influential citation count compared to their corresponding ground truth ($n=9,376$, in blue) across subperiods. a, b and c, Note that the influential citation count is retrieved from Semantic Scholar [34].
@@
Figure B6: The infuential citation bias in existing generated references is unrelated to title length, number of authors, and publication venue. This figure shows that the existing generated references ($n=9,376$, in orange) consistently exhibit a higher influential citation count compared to their corresponding ground truth ($n=9,376$, in blue) across title length, number of authors, and publication venue. a, b and c, Note that the influential citation count is retrieved from Semantic Scholar [34].
Appendix C Additional Tables
Table C1: Papers included in the analysis.
@@ Appendix C Additional Tables
Table C2: Papers excluded from the analysis.

### v3

@@ Discussion
We present an experiment to explore the intrinsic citation behavior of LLMs and their potential biases when generating scholarly references. Whereas, previous work focuses on LLMs generating short papers or literature reviews (Qureshi et al. 2023; Walters and Wilder 2023), we let GPT-4, GPT-4o and Claude 3.5 generate suggestions of scholarly references for anonymized in-text citations. Importantly, we do not enhance the LLM through search and retrieval-augmented generation, but evaluate the model’s internalization of citation patterns in its parametric knowledge obtained during training. While, our experimental setup may not fully reflect real-world usage of LLMs for citation generation, which often involves more interactivity and reliance on external data sources, it provides a controlled laboratory setting to assess the parametric knowledge and inherent biases of LLMs.
@@
Our findings are significant as they represent a first step towards understanding the real-world impact of LLMs in scientific research (Boiko, MacKnight, and Gomes 2023; Lu et al. 2024; Zheng et al. 2023). By highlighting the heightened citation bias in LLM generated references, we demonstrate the models’ tendency to favor highly cited papers, which could exacerbate existing biases in scientific discourse. This evaluation moves beyond more traditional LLM benchmarks (Srivastava et al. 2022), emphasizing the practical implications of deploying these models in academic contexts (Jimenez et al. 2024). The results suggest that while LLMs have the potential to streamline various aspects of research, careful consideration is needed to mitigate the amplification of biases, such as the “Matthew effect.”
@@
One plausible hypothesis for the heightened citation bias observed in LLMs is the increased frequency of citations to heavily cited papers within the model’s training data. This prevalence makes these references more likely to be generated accurately and recognized as existent. Additionally, such biases may stem from generic training effects, where models preferentially learn patterns that are more common in the data, leading to biases towards shorter titles, more heavily cited, and slightly less recent works (Kandpal et al. 2023). These tendencies may persist despite improvements in data quantity or model sophistication as indicated by our experiments with GPT-4o and Claude 3.5.
@@
We develop and open-source an extensible, automated pipeline to systematically analyze the references generated by LLMs. Although our methodology is robust, it is not without limitations. The use of simple prompts and the zero-shot setting (Kojima et al. 2022) aims to minimize bias in the generation process, but this simplicity might not capture the full spectrum of potential LLM capabilities. There are numerous alternative approaches and prompt designs that future research can explore to enhance the accuracy and relevance of generated references (Wang et al. 2022; Wei et al. 2022b; Yao et al. 2024). However, our iterative approach indicates that biases remain inherent in these generations. Additionally, future research can also extend the experiment beyond our specific sample of papers and observe the impact of cross-disciplinary differences in citation practices.
@@
In conclusion, while LLMs can significantly aid in citation generation, they also risk amplifying existing biases and introducing new ones, potentially skewing the structuring and the dissemination of scientific knowledge. Our study underscores the necessity for developing balanced methods to interact with LLMs, incorporating diverse datasets, and implementing bias mitigation strategies. Fair prompting techniques (Ma et al. 2023), for instance, can be employed to reduce bias, but continuous vigilance and methodological innovation are required to ensure that the integration of LLMs into academic workflows promotes accurate knowledge dissemination.
Acknowledgements
@@
Appendix A Appendix
We detail our data processing in Appendix B and show supplementary figures and tables.
Appendix B Data
We describe the steps of our automated pipeline to retrieve all the necessary information for our analysis. Our data collection resulted in $166$ papers published at AAAI ($25$), NeurIPS ($72$), ICML ($38$), and ICLR ($31$) for a total of 3,066 references (see Appendix Table LABEL:tab:paper_details for a full list of included papers). The data collection pipeline uses GPT-4-0613 to postprocess parts of the data, which costs approximately 14 dollars for our experiment. Note that these steps only have to be carried out once for the data collection. However, steps 4 and 5 are also used to postprocess and enrich the information of the generated references and will need to be carried out for each run. The experiment was run on 4 November 2023 and each step was manually verified and tested. Besides using GPT-4-0613, we also ran steps 6 and 7 for GPT-4o-2024-05-13 and Claude-3-5-sonnet-20240620 on 27 July 2024.
@@ Appendix B Data
We search for all papers on arXiv originally posted between 1 March 2022 and 31 October 2023 in the machine learning (cs.LG) category which refer to AAAI, NeurIPS, ICLR, or ICML in their journal reference. Note that we also verify whether we can use all these arXiv papers given their data licenses and attribute their participation in Appendix Table LABEL:tab:paper_details. We use keywords (i.e., workshop, tiny paper, 2020, 2021, track on datasets and benchmarks, and bridge) to remove papers that do not appear in the conference proceedings or earlier than 2022. We download and unzip the tar.gz file provided by the authors to arXiv and check whether the paper exists on Semantic Scholar via title matching. We store the title, ID, and date from arXiv and Semantic Scholar. Additionally, we store all the reference titles with their corresponding ID from Semantic Scholar (Kinney et al. 2023).
@@
Below, we share with you a list of references with their corresponding citation number between square brackets. Could you for each reference extract the authors, the number of authors, title, publication year, and publication venue? Please only return the extracted information in a markdown table with the citation number (without brackets), authors, number of authors, title, publication year, and publication venue as columns. $=$ [LLM generated reference list]
@@
We enrich the information from the introduction references by matching the extracted title from the csv file in the previous step with the reference titles that we extracted from Semantic Scholar in step 1. This approach provides an extra check that GPT-4 does not change the title information in Step 4. After matching, we can use the Semantic Scholar ID to retrieve the publication venue, year, authors, citation count, influential citation count, and reference count (Kinney et al. 2023). Additionally, we store the IDs of the papers to which the introduction references themselves refer.
@@
We prompt GPT-4-0613 with the main content, which includes the author information, conference information, abstract, and introduction, accompanied by the default system message: “You are a helpful assistant” and the following prompt:
@@
We determine whether the generated references exist via title and author names matching with Semantic Scholar entries (Kinney et al. 2023). We search Semantic Scholar for the three best matches based on the reference’s title and then compute the title and author names similarity. For titles, we measure the similarity between the Semantic Scholar match and the generated reference by comparing the best matching substring. For authors, we compare them by splitting into tokens (words), removing duplicates, and then calculating the similarity based on the best partial match of the sets of tokens. In case of “et al.,” we only consider the first author. The similarity is computed by character-level comparison. We determined the thresholds for the title and authors scores by manually labelling $100$ matches as true or false and minimizing the false positive rate. We obtain on this sample an accuracy of $95\%$ with $5$ false positives, i.e. generated references falsely classified as non-existent.
@@
We also build on our “vanilla” approach, by introducing an “iterative” approach where we prompt GPT-4-0613 with the main content accompanied by the default system message: “You are a helpful assistant” and the following prompt:
[vanilla prompt + LLM’s response ] The following references associated with these citation numbers: [numbers of non-existent generated references ] do not exist. Can you replace all these non-existent references with existing ones? Keep the other references as they are. Do not return anything except the citation number between square brackets and the corresponding reference. $=$ [main content]
@@
Figure B1: Properties of the ground truth and GPT-4 generated introduction references for the iterative strategy are consistent with the properties of the vanilla strategy. This figure displays the properties of the ground truth ($n=5,178$, in blue) and GPT-4 generated references ($n=5,178$, in green), further subdividing the generated references into existing ($n=3,244$, in orange) and non-existent categories ($n=1,934$, in red), from the original data sources of five runs for the iterative strategy with GPT-4. Note that these are the references which are labelled “non-existent” in the vanilla strategy. a, b, c, d, e, f and g, The iterative results exhibit very similar properties to the vanilla results shown in Figure 2.
@@
Figure B2: The properties of the existing GPT-4 generated references and their corresponding ground truth are consistent between the original data sources and Semantic Scholar data. This figure compares the computation of the properties of the existing GPT-4 generated references ($n=9.376$) and their corresponding ground truth ($n=9.376$) between the data from the original sources (in dark blue and orange, as shown in Figure 2) and from Semantic Scholar (in light blue and orange). a, The distributions of the number of characters in title for the existing generated references and their corresponding ground truth are very similar between the data from the original sources and Semantic Scholar. b, There is a discrepancy between the data from the original sources and Semantic Scholar for the number of authors in the existing generated references due to the extensive use of “et al”. This discrepancy results in a relatively larger portion of three authors or more, but does not change the previous conclusions. c, The distributions over time are very similar between the data from the original sources and Semantic Scholar. d, There is a discrepancy between the data from the original sources and Semantic Scholar for the publication venues. The discrepancy is consistent across the existing generated references and their corresponding ground truth as both have a lower number of arXiv papers and a larger number of ICLR papers.
@@
Figure B3: A high consistency in publication venue distributions between ground truth and GPT-4 generated references with a notable bias towards arXiv, NeurIPS, and “Others”. This figure displays the distributions and pairwise transition of publication venues for ground truth and GPT-4 generated references at the conference and individual paper and reference level. a and b, We observe that the distributions of publication venues for both ground truth and generated references are very similar across the various conferences, i.e., AAAI, NeurIPS, ICML, and ICLR. c and d, The pairwise transition matrix from ground truth to generated publication venues at the reference level indicates a large overall agreement, but with a strong preference in GPT-4 generated references for arXiv, NeurIPS, and “Others” in the case of disagreement. The preference for NeurIPS may be due to the relatively large number of NeurIPS papers in our sample and the large share of arXiv and “Others” points to favoring a wider array of venues which may potentially dilute the perceived relevance of key conferences. e, The scatter plot shows for each paper the number of ground truth references to one of the top conferences, i.e., AAAI, NeurIPS, ICML, and ICLR, and the corresponding number of generated references (×$\times 5$ for five runs) which refer to the same conference.The strong pairwise correlation between the ground truth and generated references to the top conferences at the individual paper level affirms the high consistency in publication venue distributions between ground truth and GPT-4 generated references.
Figure B4: Properties of the ground truth and GPT-4o generated introduction references are consistent with the properties of GPT-4. This figure displays the properties of the ground truth ($n=8,961$, in blue) and GPT-4o generated references ($n=8,961$, in green), further subdividing the generated references into existing ($n=6,552$, in orange) and non-existent categories ($n=2,409$, in red), from the original data sources of three runs for the vanilla strategy with GPT-4o. a, b, c, d, e, f and g, The GPT-4o generated references exhibit very similar properties to the GPT-4 results shown in Figure 2, except for the existence rate which may be due to the papers now being part of the training data and the model’s enhanced capabilities.
Figure B5: Properties of the ground truth and Claude 3.5 generated introduction references are consistent with the properties of GPT-4. This figure displays the properties of the ground truth ($n=2,893$, in blue) and Claude 3.5 generated references ($n=2,893$, in green), further subdividing the generated references into existing ($n=2,611$, in orange) and non-existent categories ($n=282$, in red), from the original data sources of three runs for the vanilla strategy with GPT-4o. a, b, c, d, e, f and g, The Claude 3.5 generated references exhibit similar properties to the GPT-4 results shown in Figure 2, except for the existence rate which may be due to the papers now being part of the training data and the model’s enhanced capabilities. Additionally, Claude 3.5, on average, generates non-existent references with shorter titles and proportionally published more in arXiv, Nature, and “others.”
Figure B6: Ground truth papers which correspond to existing GPT-4 generated references that appear in the paper have substantially more citations. This figure displays the properties of the ground truth references which correspond to existing GPT-4 generated references ($n=9,376$, in yellow), the subset of existing generated references which appear in the paper itself ($n=2,474$, in blue), and the non-existent generated references ($n=5,178$, in green), from the original data sources of five runs for the vanilla strategy with GPT-4. a, b, c, d, e and f, The ground truth papers which correspond to existing references which appear in the paper have by far the most citations, followed by the existing references, and the ground truth papers corresponding to non-existent references have the lowest numbers of citations. These findings further indicate the tendency for LLMs to more easily generate references to highly cited papers. The distributions of all other characteristics are very similar.
@@
Figure B7: The citation bias in existing GPT-4 generated references is consistent across title length, number of authors, and publication venue. This figure shows that the existing GPT-4 generated references ($n=9,376$, in orange) consistently exhibit a higher citation count compared to their corresponding ground truth ($n=9,376$, in blue) across title length, number of authors, and publication venue. a, The citation counts across the number of characters in title reveals that the discrepancy in number of citations between the existing generated and ground truth references is consistent over various title lengths. b and c, The distributions of citation counts per number of authors and publication venues show that the existing generated references consistently exhibit a higher citation count than their corresponding ground truth counterparts.
Figure B8: The influential citation bias in existing GPT-4 generated references is unrelated to the recency of ground truth references. This figure shows that the existing GPT-4 generated references ($n=9,376$, in orange) consistently exhibit a higher influential citation count compared to their corresponding ground truth ($n=9,376$, in blue) across subperiods. a, b and c, Note that the influential citation count is retrieved from Semantic Scholar (Valenzuela-Escarcega, Ha, and Etzioni 2015).
Figure B9: The infuential citation bias in existing GPT-4 generated references is unrelated to title length, number of authors, and publication venue. This figure shows that the existing GPT-4 generated references ($n=9,376$, in orange) consistently exhibit a higher influential citation count compared to their corresponding ground truth ($n=9,376$, in blue) across title length, number of authors, and publication venue. a, b and c, Note that the influential citation count is retrieved from Semantic Scholar (Valenzuela-Escarcega, Ha, and Etzioni 2015).
Vanilla Run 1 Run 2 Run 3 Run 4
Run 2 17.90
Run 3 17.11 17.30
Run 4 17.73 16.69 16.35
Run 5 18.26 17.78 17.06 18.44
Table C1: Overlap between generated sets of references of different runs by GPT-4. We see on average a $17\%$ overlap between different runs, which indicate that the models do not suffer from mode collapse (numbers in % with respect to total number of references).
Vanilla GPT-4o Claude 3.5
Run 1 Run 2 Run 3 Run 1 Run 2 Run 3
Existence 74.34 73.10 71.92 90.25 89.53 83.66
Cited in paper 32.16 33.28 33.71 42.08 41.50 39.26
Cited in introduction 24.15 26.23 25.81 34.23 33.53 31.85
Pairwise Match (PM) 10.34 10.62 10.39 18.04 17.24 14.74
PM for uniquely 17.49 19.10 18.28 29.75 29.25 25.64
Table C2: Summary statistics of generated references by GPT-4o and Claude 3.5. (numbers in % with respect to total number of references).
@@
Table C3: Papers included in the analysis.